# PowerGraph: A power grid benchmark dataset for graph neural networks

### v2

@@ Abstract
Power grids are critical infrastructures of paramount importance to modern society and, therefore, engineered to operate under diverse conditions and failures. The ongoing energy transition poses new challenges for the decision-makers and system operators. Therefore, we must develop grid analysis algorithms to ensure reliable operations. These key tools include power flow analysis and system security analysis, both needed for effective operational and strategic planning. The literature review shows a growing trend of machine learning (ML) models that perform these analyses effectively. In particular, Graph Neural Networks (GNNs) stand out in such applications because of the graph-based structure of power grids. However, there is a lack of publicly available graph datasets for training and benchmarking ML models in electrical power grid applications. First, we present PowerGraph, which comprises GNN-tailored datasets for i) power flows, ii) optimal power flows, and iii) cascading failure analyses of power grids. Second, we provide ground-truth explanations for the cascading failure analysis. Finally, we perform a complete benchmarking of GNN methods for node-level and graph-level tasks and explainability. Overall, PowerGraph is a multifaceted GNN dataset for diverse tasks that includes power flow and fault scenarios with real-world explanations, providing a valuable resource for developing improved GNN models for node-level, graph-level tasks and explainability methods in power system modeling. The dataset is available at https://figshare.com/articles/dataset/PowerGraph/22820534 and the code at https://github.com/PowerGraph-Datasets
@@ 1 Introduction
  • A dataset that contributes to the society domain within the OGB graph dataset [18].
@@
The paper is structured as follows: Section 2 outlines the dataset structure; Section 3 presents the benchmark of GNN for power flow analysis and optimal power flow analysis; Section 4 discusses the benchmark of GNN and explainability methods for cascading failure analysis; and Section 5 concludes with final remarks and discussion.
@@ 2.1 Node-level task: power flow and optimal power flow
The PowerGraph dataset for node-level tasks is obtained by performing power flow (PF) and optimal power flow (OPF) simulations with MATPOWER [20], using the formulations in Sections A.2.1 and A.2.2. This dataset consists of N attributed graphs, denoted as $\mathcal{G}={G_{1},G_{2},\ldots,G_{N}}$, with each graph representing a different grid loading condition. An attributed graph is defined as $G=(\mathcal{V},\mathcal{E},\textbf{V},\textbf{E})$, where $\mathcal{V}$ is the set of nodes (buses, i.e., the power injection points in the grid), and $\mathcal{E}$ is the set of edges (branches, i.e., the power transmission asset in the grid). The node feature matrix $\textbf{V}\in\mathbb{R}^{\lvert\mathcal{V}\rvert\times t}$ includes t features for each of the $\lvert\mathcal{V}\rvert$ nodes, and the edge feature matrix $\textbf{E}\in\mathbb{R}^{\lvert\mathcal{E}\rvert\times s}$ includes s features for each of the $\lvert\mathcal{E}\rvert$ edges. The graph connectivity information is encoded in COO format [27]. Figure 1 illustrates the dataset structure for both PF and OPF tasks. PF analysis aims to solve the power flow equations, given the power dispatched at each generator except for a balancing bus called the slack bus. The predicted node-level features depend on the type of electrical bus, as detailed in Section A.2.1. The OPF problem aims to determine the optimal generator (dispatch) to operate the grid at minimum cost while respecting physical constraints. Here too, the predicted node-level features depend on the type of bus, as detailed in Section A.2.2.
Figure 1: Instance of the PowerGraph dataset for power flow and optimal power flow. The input node features are in red, and output node-level predictions are in green. The known input quantities are reported for different node types, and the unknown quantities are set to zero (indicated by an empty cell in the picture). Similarly, the output quantities depend on the node type; if a variable is known, we mask it during training, and masked values are indicated with grey cells. The quantities are: active power generation $P_{g}$, reactive power generation $Q_{g}$, active power demand $P_{d}$, reactive power demand $Q_{d}$, voltage magnitude $V$ , and voltage angle $\theta$, the number of loads $N_{loads}$, and number of generators $N_{gen}$. The edge level features are: branch conductance $G_{ij}$ and branch susceptance $B_{ij}$.
@@ 2.2 Graph-level task: cascading failure analysis
3 Benchmarking power flow analysis: node-level tasks
@@
[…]
@@ Discussion
Table 4 shows the best models that are obtained with the GCN, GAT, GINe, and Transformer architectures for the four power systems. Full grid search results are available at powergraph.ivia.ch. GAT is generally the best-performing model. Notably, GCN disregards edge features. This, however, does not significantly impact the performance of node-level tasks. We observe that the best performance is consistently achieved with a single MPL for all cases. As expected, the OPF task is more complex, leading to a higher mean squared error (MSE) compared to the PF task. Figure 3 details the node-averaged MSE for each predicted physical quantity. Although the global MSE is consistently lower for the PF task, Figure 3 reveals that the PF models do not necessarily outperform the OPF models in predicting voltage magnitude and angle, particularly for the IEEE118-bus system. In the PF problem, the generated power is known except for the slack bus, resulting in more accurate predictions for $P_{g}$ and $Q_{g}$. However, the models on the IEEE39 and IEEE118 grids exhibit the worst performance on both PF and OPF tasks. Conversely, the PF model performs best on the UK grid, but the MSE increases in the OPF task, while the IEEE24 grid models exhibit strong performance in both tasks. The IEEE24 and UK grids have fewer nodes compared to the other two systems, simplifying node-level regression tasks.
@@
4 Benchmarking cascading failure analysis - graph-level tasks: classification, regression, and explanability
Experimental setting and evaluation metrics
We utilize baseline GNN architectures commonly used in the graph xAI community for each power grid dataset. Specifically, we use GCNConv [28], GATConv [29], GINEConv [30] and TransformerConv [31] to demonstrate that the PowerGraph datasets can benchmark GNNs and their explanatory methods. Additionally, we experiment with state-of-the-art graph transformer convolutional layers [31], which form the backbone of recent Graph Transformer models such as GraphGPS [33], Transformer-M [34], and TokenGT [35]. These models are chosen because they incorporate edge features, which are highly relevant for power grids. The GNN architecture includes multiple MPLs, each followed by a PReLU activation function [32], a maximum global pooling operator to obtain graph-level embeddings from node embeddings, and one fully connected layer. The Adam optimizer is used with an initial learning rate of $10^{-3}$. We tune the number of MPLs, selecting from 1, 2, 3, and the hidden dimensionality, selecting from 8, 16, 32. Each model is trained for 200 epochs, with the learning rate adjusted automatically using a scheduler when a metric stops improving. The datasets are split into 85% train, 5% validation, and 10% test sets, with a batch size of 16. For classification models, balanced accuracy [36] is used as the evaluation metric due to the strong class imbalance (see Table 3). For regression models, mean squared error is used as the reference metric.
@@
Discussion
We report the best model performance over the GCN, GAT, GINe, and Transformer models and different combinations of hyperparameters for the four power systems and MPL in Tables 5. Full grid search results and additional performance metrics are available at powergraph.ivia.ch. Among these, only GCN does not consider edge features, resulting in lower performance in most cases. Transformer and GINe are the best-performing models. Overall, the binary and multi-class classification models exhibit good results. However, the regression model, which predicts the exact demand not served, does not achieve the desired level of performance. Indeed, the R2 score [37], also known as the coefficient of determination, achieves a maximum value of 0.43 for the best regression model on the IEEE24 dataset. However, the R [37] score for the remaining power systems remains below 0.26. Ideally, we aim for an R score closer to 1 to indicate a better fit for the model. It is evident that future research should focus on developing approaches to improve the performance of GNN on regression tasks. In parallel, even small improvements in classification tasks should not be underestimated, as they can greatly influence decision-making in critical infrastructure. Thus, it is essential to have an exceptionally precise and accurate model to ensure that system operators can safely operate the power grid.
4.1 Benchmarking explanations on the graph-classification models
@@
The PowerGraph benchmark with explanations tests and compares explainability methods. The role of explainers is to identify the edges that are necessary for the graphs to be classified as Category A. This choice is explained in Appendix A.4. Then, the resulting edges are evaluated on how well they match the explanation masks, which represent the cascading edges, using the accuracy score, and on how they contribute to the model’s prediction, using the faithfulness metric. The accuracy score refers to the balanced accuracy metric, defined in Appendix A.5, between the ground truth target cascading failure edges and the estimated explanatory edges by the xAI method. The faithfulness score measures the changes in model outputs induced by retraining only with the important graph entities identified, and its definition is given in Appendix B.5. We compare the results obtained for each of the four PowerGraph datasets: IEEE24, IEEE39, IEEE118, and UK. For each dataset, we select the trained Transformer model with 3 layers and 16 hidden units, obtained with the configuration of Section 4. To benchmark explainability methods, having the best GNN model is not essential. By appropriately filtering predictions (correct or mixed) and focusing the explanations (on the phenomenon or model) [38], we can work around lower test accuracy. We compare non-generative methods, including the heuristic Occlusion [39], gradient-based methods Saliency [40], Integrated Gradient [41], and Grad-CAM [42], and perturbation-based methods GNNExplainer [43], PGMExplainer [44] and SubgraphX [45]. We also consider generative methods: GSAT [46], D4Explainer [47] and RCExplainer [48]. For more details on generative vs. non-generative explainers, see Appendix B.2 and B.3. We compare those explainability methods to the base estimators: Random, Truth, and Inverse. Random assigns random importance to edges following a uniform distribution. Truth estimates edge importance as the pre-defined ground-truth explanations of the datasets. i.e., the cascading edges. The Inverse estimator represents the worst-case scenario, assigning edges opposite to the ground truth. Appendix B provides additional experimental details on Transformer performance and the explainability methods.
@@ 5 Conclusions
PowerGraph addresses critical challenges encountered in utilizing Graph Neural Networks (GNNs) for power systems analyses. PowerGraph effectively fills a significant gap in the GNN domain for power engineering applications by offering a dataset tailored for node and graph-level tasks and model explainability. Through rigorous benchmarking against a range of GNN and explainability models, PowerGraph exhibits high performance in graph classification, albeit indicating a need for further refinement in regression models. Regression models remain essential in power systems and can be solved by GNNs for tasks like power flow analysis and fault scenarios. However, GNN still encounters challenges in this context. Additionally, the complexity of architecture varies considerably between node-level and graph-level tasks. For instance, power flow analysis at the node level requires fewer message-passing layers compared to cascading failure analysis at the graph level. In power flow analysis, we conclude that nearest neighbor information is crucial for accurate output prediction. However, this is not the case for predicting cascading failures, which requires information from more distant nodes for accurate predictions. PowerGraph is the first real-world dataset to offer ground-truth explanations for graph-level tasks in explainable AI, enabling thorough evaluation of various explainability methods. Despite this advancement, the performance of these methods remains suboptimal, highlighting the lack of a universally superior approach. Given the crucial role of explainability for power grid operators, this underscores the ongoing need for dedicated research and development in this field.
@@
      Did you describe the limitations of your work? [Yes] See: Section 3 paragraph ’Discussion’, Section 4 paragraph ’Discussion’, Section 4.1 paragraph ’Discussion’ and the Section 5.
@@
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Sections 3, 4 and 4.1, and the respective paragraphs ’Experimental setting and evaluation metrics’
@@
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See Table 4, Table 5 andSection 4.1 paragraph ’Results’.

### v3

@@ Abstract
Power grids are critical infrastructures of paramount importance to modern society and, therefore, engineered to operate under diverse conditions and failures. The ongoing energy transition poses new challenges for the decision-makers and system operators. Therefore, developing grid analysis algorithms is important for supporting reliable operations. These key tools include power flow analysis and system security analysis, both needed for effective operational and strategic planning. The literature review shows a growing trend of machine learning (ML) models that perform these analyses effectively. In particular, Graph Neural Networks (GNNs) stand out in such applications because of the graph-based structure of power grids. However, there is a lack of publicly available graph datasets for training and benchmarking ML models in electrical power grid applications. First, we present PowerGraph, which comprises GNN-tailored datasets for i) power flows, ii) optimal power flows, and iii) cascading failure analyses of power grids. Second, we provide ground-truth explanations for the cascading failure analysis. Finally, we perform a complete benchmarking of GNN methods for node-level and graph-level tasks and explainability. Overall, PowerGraph is a multifaceted GNN dataset for diverse tasks that includes power flow and fault scenarios with real-world explanations, providing a valuable resource for developing improved GNN models for node-level, graph-level tasks and explainability methods in power system modeling. The dataset is available at https://figshare.com/articles/dataset/PowerGraph/22820534 and the code at https://github.com/PowerGraph-Datasets
@@ 1 Introduction
  • A dataset that contributes to the society domain within the OGB graph dataset taxonomy [18].
@@
The paper is structured as follows: Section 2 outlines the dataset structure; Section 3 presents the benchmark of GNN for power flow analysis, optimal power flow and cascading failure analysis, Section 4 presents explainability methods for cascading failure analysis; and Section 6 concludes with final remarks and discussion.
@@ 2.1 Node-level task: power flow and optimal power flow
The PowerGraph dataset for node-level tasks is obtained by performing power flow (PF) and optimal power flow (OPF) simulations with MATPOWER [20], using the formulations in Sections A.2.1 and A.2.2. This dataset consists of N attributed graphs, denoted as $\mathcal{G}={G_{1},G_{2},\ldots,G_{N}}$, with each graph representing a different grid loading condition. An attributed graph is defined as $G=(\mathcal{V},\mathcal{E},\textbf{V},\textbf{E})$, where $\mathcal{V}$ is the set of nodes (buses, i.e., the power injection points in the grid), and $\mathcal{E}$ is the set of edges (branches, i.e., the power transmission asset in the grid). The node feature matrix $\textbf{V}\in\mathbb{R}^{\lvert\mathcal{V}\rvert\times t}$ includes t features for each of the $\lvert\mathcal{V}\rvert$ nodes, and the edge feature matrix $\textbf{E}\in\mathbb{R}^{\lvert\mathcal{E}\rvert\times s}$ includes s features for each of the $\lvert\mathcal{E}\rvert$ edges. Figure 1 illustrates the dataset structure for both PF and OPF tasks. PF analysis aims to solve the power flow equations, given the power dispatched at each generator except for a balancing bus called the slack bus. The predicted node-level features depend on the type of electrical bus, as detailed in Section A.2.1. The OPF problem aims to determine the optimal generator (dispatch) to operate the grid at minimum cost while respecting physical constraints. Here, too, the predicted node-level features depend on the type of bus, as detailed in Section A.2.2.
Figure 1: Instance of the PowerGraph dataset for power flow and optimal power flow. The input node features are in red, and output node-level predictions are in green. The known input quantities are reported for different node types, and the unknown quantities are set to zero to maintain the dataset’s dimensionality structure (indicated by an empty cell in the picture). Similarly, the output quantities depend on the node type; if a variable is known, we mask it during training, and masked values are indicated with grey cells. The quantities are: active power generation $P_{g}$, reactive power generation $Q_{g}$, active power demand $P_{d}$, reactive power demand $Q_{d}$, voltage magnitude $V$ , and voltage angle $\theta$, the number of loads $N_{loads}$, and number of generators $N_{gen}$. The edge level features are: branch conductance $G_{ij}$ and branch susceptance $B_{ij}$.
@@ 2.2 Graph-level task: cascading failure analysis
3 Benchmarking the PowerGraph dataset
In this section, we present the experimental setting to benchmark the PowerGraph dataset for node-level tasks (i.e. power flow and optimal power flow analysis) and graph level tasks (i.e. cascading failure analysis).
@@ 3 Benchmarking the PowerGraph dataset
For each power grid dataset, both for graph and node-level tasks we utilize four baseline GNN architectures: GCNConv [27], GATConv [28], GINEConv [29], and TransformerConv [30], which are commonly used in the graph xAI community. We experiment with state-of-the-art graph transformer convolutional layers [30], forming the backbone of recent models such as GraphGPS [31], Transformer-M [32], and TokenGT [33], chosen for their ability to incorporate edge features relevant to power grids. The GNN architecture includes multiple message passing layers (MPLs), each followed by a PReLU activation function [34]. When graph-level tasks are involved, a maximum global pooling operator for obtaining graph-level embeddings from node embeddings, and one fully connected layer. We perform a grid search over the number of MPLs (1, 2, 3) and the hidden dimensionality (8, 16, 32). The Adam optimizer is used with an initial learning rate of $10^{-3}$, and each model is trained for 200 epochs, with the learning rate adjusted using a scheduler that automatically reduces it if a metric stops improving. The datasets are split into 85% train, 5% validation, and 10% test sets, with a batch size of 32 for node-level tasks and 16 for graph-level tasks. For classification models, balanced accuracy [35] is used due to class imbalance, (see Table 3). For regression models, mean squared error is used as the reference metric.
Observations
Table 4 presents the best models for node-level tasks using GCN, GAT, GINe, and Transformer architectures across four power systems. Similarly, Table 5 reports the top-performing models for graph-level tasks with the same architectures and various hyperparameter combinations. In addition to the presented result in this paper and its appendices, we present the complete results, including all models and their performances, on our website: https://powergraph.ivia.ch. The website also includes Gradient Boosted Trees results as a baseline for comparison. Figure 3 illustrates the node-averaged MSE for each predicted physical quantity.
@@ Discussion
We train our models using data from current simulation methods (see A.2.1 and A.3), aiming for the highest accuracy in classification and regression to replicate those results, while the key advantage of our ML models lies in surpassing traditional solvers or simulations in computational efficiency. Concerning node-level tasks, the OPF task is inherently more complex, resulting in a higher mean squared error (MSE) for the ML model trained on it compared to the one trained on the PF task, as illustrated in Figure 3, which presents the node-averaged MSE for each predicted physical quantity. While the global MSE is consistently lower for the PF task, the PF models do not always outperform the OPF models in predicting voltage magnitude and angle, particularly for the IEEE118-bus system. In the PF problem, where generated power is known except for the slack bus, predictions for $P_{g}$ and $Q_{g}$ are more accurate. However, the models on the IEEE39 and IEEE118 grids show the worst performance across both tasks. Conversely, the PF model performs best on the UK grid, though the MSE increases in the OPF task, whereas the IEEE24 grid models show strong performance in both tasks. The fewer nodes in the IEEE24 and UK grids simplify node-level regression tasks.
Notably, for node-level tasks, the GAT model generally outperforms others. Nevertheless, the GCN model, which disregards edge features, also shows strong results without significantly impacting performance. The best results are consistently achieved with a single MPL across all cases. For graph-level tasks, the GCN model consistently shows lower performance compared to other models. This highlights a clear distinction from power flow tasks, where edge features like line limits are crucial for modeling potential cascades and therefore have a significant impact on performance. In contrast, the Transformer and GINe models emerge as the best performers. The Transformer model leverages its attention mechanism to dynamically weigh the significance of neighboring nodes, capturing complex relationships and leading to superior learning outcomes. This attention-based approach is a key factor in the Transformer’s and GAT’s effectiveness across both node-level and graph-level tasks. Additionally, the GINe model demonstrates robust performance due to its strong capability in capturing and aggregating local neighborhood information, which enhances its predictive accuracy [29].
While binary and multi-class classification models show good results, the regression model predicting the exact demand not served does not. The R2 [36] score, also known as the coefficient of determination, reaches a maximum of 0.43 for the best regression model on the IEEE24 dataset, but remains below 0.26 for other power systems. Ideally, an R2 score closer to 1 is desired for a better model fit. This highlights the need for future research to focus on improving GNN performance on regression tasks both at the graph and node-level. We observe the limitations of GNNs, particularly in node regression tasks, where simpler models like Gradient Boosted Trees (GBT) outperform GNNs (see https://powergraph.ivia.ch). However, this trend does not extend to graph-level tasks, where GNNs consistently prove to be the superior method. This highlights the need for further development of GNN models tailored specifically for node regression tasks, which are critical in power systems analysis. At the same time, even minor improvements in classification tasks should not be underestimated, as they can significantly impact decision-making in critical infrastructures, as power grids.
@@
4 Benchmarking explanations on the graph-classification models
@@
The PowerGraph benchmark with explanations tests and compares explainability methods. Explanations are subsets of the grids. We prefer generating graph-based explanations over textual ones, as graph structures are more intuitive and effective for conveying critical nodes and edges to experts in transmission grids. The role of explainers is to identify the edges that are necessary for the graphs to be classified as Category A.This choice is explained in Appendix A.4. Then, the resulting edges are evaluated on how well they match the explanation masks, which represent the cascading edges, using the accuracy score, and on how they contribute to the model’s prediction, using the faithfulness metric. The accuracy score refers to the balanced accuracy metric, defined in Appendix A.6, between the ground truth target cascading failure edges and the estimated explanatory edges by the xAI method. The faithfulness score measures the changes in model outputs induced by retraining only with the important graph entities identified, and its definition is given in Appendix B.5. We compare the results obtained for each of the four PowerGraph datasets: IEEE24, IEEE39, IEEE118, and UK. For each dataset, we select the trained Transformer model with 3 layers and 16 hidden units, obtained with the configuration of Section 3. To benchmark explainability methods, having the best GNN model is not essential. By appropriately filtering predictions (correct or mixed) and focusing the explanations (on the phenomenon or model) [37], we can work around lower test accuracy. We compare non-generative methods, including the heuristic Occlusion [38], gradient-based methods Saliency [39], Integrated Gradient [40], and Grad-CAM [41], and perturbation-based methods GNNExplainer [42], PGMExplainer [43] and SubgraphX [44]. We also consider generative methods: GSAT [45], D4Explainer [46] and RCExplainer [47]. For more details on generative vs. non-generative explainers, see Appendix B.2 and B.3. We compare those explainability methods to the base estimators: Random, Truth, and Inverse. Random assigns random importance to edges following a uniform distribution. Truth estimates edge importance as the pre-defined ground-truth explanations of the datasets. i.e., the cascading edges. The Inverse estimator represents the worst-case scenario, assigning edges opposite to the ground truth. Appendix B provides additional experimental details on Transformer performance and the explainability methods.
@@ Discussion
5 Limitations
While this study provides a valuable benchmark for power grid analysis using GNN models, there are limitations that should be acknowledged. Using power grid models, such as the IEEE test case systems, is a standard practice in power engineering modeling and analyses. Although these power systems are not ideal, they are still based on real power grids. The main reason we utilize these grids in our study is because real power system data is safety-critical and often classified, making it inaccessible for research. Furthermore, cascading failures are rare events, and data on their progression and the conditions under which they occur are largely unavailable. As a result, synthetic data is the only viable option for cascading failure analysis, which could limit the generalizability of the findings to real-world scenarios. Nevertheless, we use the Cascades model to simulate cascading failures and generate GNN datasets, which have been validated with historical blackout data from the WECC grid [48]. Furthermore, the small number of buses in the IEEE test cases may constrain the ability of the models to capture long-range dependencies, potentially affecting their performance on larger, more complex power grids. While this work demonstrates a method for modeling power systems with GNNs, other approaches could be explored in future research to enhance model accuracy and robustness. For instance, incorporating energy price variations to analyze the multi-period economic dispatch, multi-period OPF and unit commitment could provide a more comprehensive understanding of the economic aspects of power grid operations, which was not considered in this study
@@ 6 Conclusions
PowerGraph addresses critical challenges encountered in utilizing Graph Neural Networks (GNNs) for power systems analyses. PowerGraph effectively fills a significant gap in the GNN domain for power engineering applications by offering a dataset tailored for node and graph-level tasks and model explainability. Through rigorous benchmarking against a range of GNN and explainability models, PowerGraph exhibits high performance in graph classification, albeit indicating a need for further refinement in regression models. Regression models remain essential in power systems and can be solved by GNNs for tasks like power flow and system security analysis. However, GNN still encounters challenges that demand further research and development. Additionally, the complexity of architecture varies considerably between node-level and graph-level tasks. For instance, power flow analysis at the node level requires fewer message-passing layers compared to cascading failure analysis at the graph level. In power flow analysis, we conclude that nearest neighbor information is crucial for accurate output prediction. However, this is not the case for predicting cascading failures, which requires information from more distant nodes for accurate predictions. PowerGraph is the first real-world dataset to offer ground-truth explanations for graph-level tasks in explainable AI, enabling thorough evaluation of various explainability methods. Despite this advancement, the performance of these methods remains suboptimal, highlighting the lack of a universally superior approach. Given the crucial role of explainability for power grid operators, this underscores the ongoing need for dedicated research and development in this field. Looking forward, we plan to enhance PowerGraph by adding a temporal graph dataset to facilitate in-depth analysis of the stages of cascading failures. Although this information is already present in the dataset, please refer to A.5, it was beyond the scope of this work. We will also benchmark methods on larger synthetic power systems, such as the grid in [49], to assess whether deeper GNN architectures are necessary, as noted in [50]. A raw dataset is available at https://powergraph.ivia.ch.
@@
      Did you describe the limitations of your work? [Yes] See: Section 3 paragraph ’Discussion’, Section 4 paragraph ’Discussion’ and the Section 5.
@@
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Sections 3 and 4, and the respective paragraphs ’Experimental setting and evaluation metrics’
@@
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See Table 4, Table 5 andSection 4 paragraph ’Results’.